# reference.bib

### v1

@@ 4.4 Ablation Study
We conduct systematic ablations on the GRAB [GRAB:2020] dataset to validate the necessity of our core components through controlled experiments. Each component is carefully analyzed to demonstrate its contribution to robust hand-object interaction generation, the details can be found in Table 4.
@@
Table 4: Ablation Study
@@
Ours 51.34$\pm$0.85 0.27$\pm$0.01 28.29$\pm$0.62 3.61$\pm$0.09 19.91$\pm$0.63
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (62406195, 62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
[…]
@@ Appendix A Implementation Details
We initialize our model from the ShapeLLM-7B checkpoint, freezing its 3D encoder and augmenting the visual backbone with Uni3D for robust dense 3D prediction, while the projection head is implemented as a shallow MLP, and LoRA is applied to streamline fine-tuning. Training unfolds in two stages: first, we optimize for seven epochs with AdamW (learning rate $2\cdot 10^{-4}$, zero weight decay) under a cosine-annealing schedule and a 2% linear warm-up; then, we continue for three additional epochs with AdamW (learning rate $5\cdot 10^{-4}$, zero weight decay) under the same cosine schedule but a 1% warm-up.
@@
We employ a T = 1000-step noising process with a cosine noise schedule, and inject positional information at both the frame- and agent-levels using sinusoidal encodings. During sampling, we apply classifier-free guidance by randomly substituting 10% of conditioning inputs with unconditional noise while retaining 90% of the original conditions, and use a guidance scale of 2.5 to steer the denoising trajectory.
@@ B.1 Affordance Reasoning
We take as input a point cloud of an object, sampled to $N$ points. The backbone is a ReCon++ [qi2024shapellm] network (or a similar architecture) that processes these points and produces per-point feature representations
@@
Multi-Token Fusion Mechanism.
@@
which is concatenated with the scale-specific image features and fed into the pixel decoder to produce the final segmentation mask. Experiments on the MUSE validation set indicate that increasing $N$ from 1 to 3 improves cIoU, demonstrating that the multi-token fusion mechanism captures more nuanced semantic details and significantly enhances fine-grained segmentation.
@@ B.2 Instruction Decomposition
OpenHOI decomposes a single high-level instruction into an ordered sequence of actionable affordance steps. Each step is marked by a special token and then grounded spatially in the 3D point cloud.
@@
We take a natural-language instruction $T_{\mathrm{ins}}$ as the model input. The backbone is a LLaMA-style Transformer that produces token-wise hidden states $\{h_{t}\}_{t=1}^{T}$ and aggregates them via a pooling operation into a single embedding
@@
Extend the MLLM’s vocabulary by adding a special marker <AFF>, which explicitly denotes the boundary of each sub-task in the generated sequence.
@@
[…]
@@
Boundary Localization & Hidden-State Extraction.
Record the positions $t_{1},\dots,t_{S}$ where <AFF> appears. For each $i=1,\dots,S$, extract the corresponding last-layer hidden vector
@@
which encodes the full context immediately preceding the end of sub-task $i$.
@@
For each step $i$, use the query $E_{i}$ to perform cross-attention over the point features $F_{\mathrm{obj}}$ and decode a per-point mask:
@@
where $\widetilde{F}$ are the fused features and $\sigma$ is the sigmoid activation. Collect the ordered set $\{M_{1},M_{2},\dots,M_{S}\}$ to obtain the final sequence of affordance masks, each aligned with its corresponding sub-task.
@@ B.3 Diffusion Process
[…]
@@
$L_{\text{hoi\_diff}}(\hat{X}_{\theta},\textbf{C})=\mathbb{E}_{X_{0}\sim p(X_{0}),X_{t}\sim p(X_{t}|X_{0}),t\sim[1,T]}||X_{0}-X_{\theta}(X_{t},t,\textbf{C})||_{2}^{2}$ (21)
@@ Appendix C Additional Experiments
C.1 Results on Extreme-Case: Completely Unseen Datasets
@@
We further subject our model to extreme-case testing to stress its generalization under the most challenging conditions. All objects and instructions in the H2O dataset are entirely novel to models trained on GRAB and ARCTIC, making evaluation on H2O a particularly stringent test of generalization. Despite this extreme distribution shift, our experiments demonstrate that the proposed model nonetheless delivers robust and state-of-the-art performance (shown in Table A1).
@@
MDM[tevet2023human] 95.12$\pm$3.21 0.64$\pm$0.04 70.54$\pm$2.75 2.36$\pm$0.11 12.50$\pm$0.50
TM2T[chuan2022tm2t] 90.45$\pm$4.50 0.68$\pm$0.03 65.47$\pm$3.20 2.51$\pm$0.10 14.45$\pm$0.60
MotionGPT[jiang2024motiongpt] 85.38$\pm$4.00 0.61$\pm$0.02 60.12$\pm$2.80 2.73$\pm$0.13 15.93$\pm$0.70
Text2HOI[cha2024text2hoi] 80.25$\pm$3.80 0.63$\pm$0.025 55.23$\pm$2.50 1.90$\pm$0.14 17.00$\pm$0.80
Ours 75.78$\pm$4.68 0.52$\pm$0.49 51.33$\pm$3.41 3.07$\pm$0.27 18.15$\pm$1.48
ARCTIC MDM[tevet2023human] 105.32$\pm$5.00 0.80$\pm$0.04 75.89$\pm$3.50 1.90$\pm$0.09 11.27$\pm$0.35
TM2T[chuan2022tm2t] 98.47$\pm$4.80 0.82$\pm$0.03 72.55$\pm$3.30 2.10$\pm$0.10 13.05$\pm$0.45
MotionGPT[jiang2024motiongpt] 93.15$\pm$4.50 0.74$\pm$0.02 65.37$\pm$3.00 2.30$\pm$0.11 12.96$\pm$0.55
Text2HOI[cha2024text2hoi] 88.02$\pm$4.30 0.71$\pm$0.025 60.28$\pm$2.80 1.50$\pm$0.12 14.78$\pm$0.65
Ours 81.36$\pm$5.77 0.63$\pm$0.12 55.78$\pm$3.62 2.69$\pm$0.43 15.44$\pm$1.48
C.2 3D MLLM Fine-tuning
We fine-tune the MLLM on the Affordance dataset[yu2025seqaffordsequential3daffordance] and the HOI dataset[GRAB:2020, fan2023arctic], We first perform coarse-grained fine-tuning on the Affordance dataset to instill strong affordance priors, and then carry out fine-grained tuning on the HOI dataset to produce our final model.The results shown in Table. A2.
@@ C.3 Sensitivity Analysis
0.5 52.23$\pm$1.06 0.40$\pm$0.01 31.05$\pm$1.56 1.97$\pm$0.19 12.77$\pm$0.45
2.0 46.04$\pm$1.19 0.28$\pm$0.03 20.98$\pm$2.30 2.62$\pm$0.04 14.96$\pm$0.61
2.5 45.15$\pm$0.94 0.25$\pm$0.04 19.74$\pm$0.16 2.65$\pm$0.03 15.25$\pm$1.44
3.0 46.55$\pm$1.74 0.27$\pm$0.02 21.03$\pm$0.67 2.68$\pm$0.15 15.03$\pm$1.75
5.0 53.25$\pm$2.07 0.38$\pm$0.02 32.85$\pm$2.04 3.40$\pm$0.06 12.86$\pm$2.07
Unseen 0.5 51.67$\pm$1.14 0.35$\pm$0.02 31.54$\pm$0.68 2.07$\pm$0.05 10.18$\pm$0.26
2.0 47.70$\pm$0.88 0.30$\pm$0.02 20.81$\pm$2.12 2.46$\pm$0.07 12.36$\pm$0.89
2.5 Ours 47.25$\pm$0.39 0.28$\pm$0.03 20.05$\pm$0.80 2.49$\pm$0.08 12.66$\pm$0.71
3.0 47.76$\pm$0.66 0.29$\pm$0.01 21.07$\pm$0.41 2.51$\pm$0.28 12.50$\pm$1.77
5.0 54.19$\pm$0.89 0.34$\pm$0.02 27.32$\pm$0.56 2.21$\pm$0.06 10.53$\pm$0.69
C.4 Ablation Study on ARCTIC
We also conducted ablation experiments on the ARCTIC dataset, and the results TableA5 showed that each of our modules played a role in generating HOI sequences
Table A5: Ablation Study on ARCTIC
@@ C.4 Ablation Study on ARCTIC
Seen GT - - - 3.39
Ours 45.15$\pm$0.94 0.25$\pm$0.04 19.74$\pm$0.35 2.65$\pm$0.16 15.25$\pm$1.44
Ours 47.25$\pm$0.39 0.28$\pm$0.03 20.05$\pm$0.80 2.49$\pm$0.08 12.66$\pm$0.71
C.5 Ablation Study on Motion In-between Refinement
@@
To the best of our knowledge, no previous work has defined an evaluation metric for motion in-between within hand-object interaction. We thus introduce a simple yet effective measure, the “Smooth Rate,” to quantify the temporal continuity of interpolated motion segments as follows,
@@
C.6 Qualitative results Compare with SOTA
@@ Appendix D Code and Dataset
We will release our code and dataset once the paper is accepted.
D.1 Dataset
@@
This dataset contains 571,645 synchronized multi-view RGB-D frames captured with five Kinect sensors in three indoor scenes. Each frame includes 3D poses for both hands, 6-DoF object poses, and verb–noun action labels (36 classes). Split into training (344,645), validation (73,380), and test (153,620) frames, H2O supports egocentric interaction recognition and manipulation benchmarks.

### v2

@@ 4.4 Ablation Study
We conduct systematic ablations on the GRAB [GRAB:2020] and ARCTIC [fan2023arctic] dataset to validate the necessity of our core components through controlled experiments. Each component is carefully analyzed to demonstrate its contribution to robust hand-object interaction generation, the details can be found in Table 4 and Table 5.
@@
Table 4: Ablation Study on GRAB
@@
[…]
Table 5: Ablation Study on ARCTIC
[…]
Seen GT - - - 3.39
[…]
4.5 Failure cases and error analysis
Given a scenario, there is a row of cabinets, and the instruction "Open the cabinet", our model can open one but cannot target a specific one (e.g., "Open the second cabinet"). This limitation arises because the model has not been trained on a large-scale 3D QA dataset, resulting in reduced logical reasoning capabilities. Due to the accumulation of model errors, performance typically degrades after more than three consecutive actions(more than 450 frames).
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (62406195, 62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), and Shanghai Engineering Research Center of Intelligent Vision and Imaging. This work was also supported in part by computational resources provided by Fcloud CO., LTD.
@@ Appendix A Implementation Details
We initialize our model from the ShapeLLM‑7B checkpoint, freezing its 3D encoder and augmenting the visual backbone with Uni3D for robust dense 3D prediction, while the projection head is implemented as a shallow MLP, and LoRA is applied to streamline fine‑tuning. Training unfolds in two stages: first, we optimize for seven epochs with AdamW (learning rate $2\cdot 10^{-4}$, zero weight decay) under a cosine‑annealing schedule and a 2% linear warm‑up; then, we continue for three additional epochs with AdamW (learning rate $5\cdot 10^{-4}$, zero weight decay) under the same cosine schedule but a 1% warm‑up.
@@
We employ a T = 1000‑step noising process with a cosine noise schedule, and inject positional information at both the frame‑ and agent‑levels using sinusoidal encodings. During sampling, we apply classifier‑free guidance by randomly substituting 10% of conditioning inputs with unconditional noise while retaining 90% of the original conditions, and use a guidance scale of 2.5 to steer the denoising trajectory.
@@ B.1 Affordance Reasoning
We take as input a point cloud of an object, sampled to $N$ points. The backbone is a ReCon++ [qi2024shapellm] network (or a similar architecture) that processes these points and produces per‑point feature representations
@@
Multi‑Token Fusion Mechanism.
@@
which is concatenated with the scale‑specific image features and fed into the pixel decoder to produce the final segmentation mask. Experiments on the MUSE validation set indicate that increasing $N$ from 1 to 3 improves cIoU, demonstrating that the multi‑token fusion mechanism captures more nuanced semantic details and significantly enhances fine‑grained segmentation.
@@ B.2 Instruction Decomposition
OpenHOI decomposes a single high‑level instruction into an ordered sequence of actionable affordance steps. Each step is marked by a special token and then grounded spatially in the 3D point cloud.
@@
We take a natural‐language instruction $T_{\mathrm{ins}}$ as the model input. The backbone is a LLaMA‐style Transformer that produces token‐wise hidden states $\{h_{t}\}_{t=1}^{T}$ and aggregates them via a pooling operation into a single embedding
@@
Extend the MLLM’s vocabulary by adding a special marker <AFF>, which explicitly denotes the boundary of each sub‑task in the generated sequence.
@@
$\texttt{Pick}\;\to\;\texttt{<AFF>}\quad\texttt{Twist}\;\to\;\texttt{<AFF>}\quad\texttt{Lift}\;\to\;\texttt{<AFF>}$
@@
Boundary Localization & Hidden‑State Extraction.
Record the positions $t_{1},\dots,t_{S}$ where <AFF> appears. For each $i=1,\dots,S$, extract the corresponding last‑layer hidden vector
@@
which encodes the full context immediately preceding the end of sub‑task $i$.
@@
For each step $i$, use the query $E_{i}$ to perform cross‑attention over the point features $F_{\mathrm{obj}}$ and decode a per‑point mask:
@@
where $\widetilde{F}$ are the fused features and $\sigma$ is the sigmoid activation. Collect the ordered set $\{M_{1},M_{2},\dots,M_{S}\}$ to obtain the final sequence of affordance masks, each aligned with its corresponding sub‑task.
@@ B.3 Diffusion Process
$\log p_{\theta}(X_{0}|\textbf{C})$ $=\log\int p_{\theta}(X_{0:T}|\textbf{C})dX_{1:T}$ (16)
$=\log\int\frac{p_{\theta}(X_{0:T}|\textbf{C})p(X_{1:T}|X_{0},\textbf{C})}{p(X_{1:T}|X_{0},\textbf{C})}dX_{1:T}$ (17)
@@
[…]
@@ Appendix C Additional Experiments
C.1 Results on Extreme‑Case: Completely Unseen Datasets
@@
We further subject our model to extreme‑case testing to stress its generalization under the most challenging conditions. All objects and instructions in the H2O dataset are entirely novel to models trained on GRAB and ARCTIC, making evaluation on H2O a particularly stringent test of generalization. Despite this extreme distribution shift, our experiments demonstrate that the proposed model nonetheless delivers robust and state‑of‑the‑art performance (shown in Table A1).
@@
[…]
C.2 Evaluation for physical realism
We supplemented our experiments by evaluating Physical Realism and IV metrics against the closest baseline, Text2HOI (HOIGPT’s code is not publicly available) in A2, and conducted ablation studies on our Physical Refinement module in A3.
Table A2: Comparison with Text2HOI
Method Physical realism $\uparrow$ IV $\downarrow$
Seen
[…]
Unseen
[…]
Table A3: Ablation Study on Physical Refinement
Method Physical realism $\uparrow$ IV $\downarrow$
Seen
[…]
Unseen
[…]
C.3 3D MLLM Fine-tuning
We fine-tune the MLLM on the Affordance dataset [yu2025seqaffordsequential3daffordance] and the HOI dataset [GRAB:2020, fan2023arctic], We first perform coarse‑grained fine‑tuning on the Affordance dataset to instill strong affordance priors, and then carry out fine‑grained tuning on the HOI dataset to produce our final model [wang2025sdevalsafetydynamicevaluation]. The results shown in Table. A4.
@@ C.4 Sensitivity Analysis
[…]
C.5 Ablation Study on Multi <AFF>
The additional ablation study results are as follows A7.
Table A7: Ablation on Affordance Configuration (Single vs. Multi)
Method MPJPE $\downarrow$ FOL $\downarrow$ FID $\downarrow$ Diversity $\rightarrow$ MModality $\uparrow$
Seen
[…]
Unseen
[…]
C.6 Ablation Study on Motion In-between Refinement
@@
To the best of our knowledge, no previous work has defined an evaluation metric for motion in‑between hand‑object interaction. We thus introduce a simple yet effective measure, the “Smooth Rate,” to quantify the temporal continuity of interpolated motion segments as follows,
@@
To determine the most suitable window size [barquero2024seamless] for the motion in-between algorithm, we conducted the following comparative experimentsA10.
Table A10: Window Size Compare on GRAB
Size SmoothRate (Seen) $\downarrow$ SmoothRate (Unseen) $\downarrow$
1 $4.77\pm 0.68$ $5.26\pm 2.47$
3 $3.59\pm 0.81$ $4.58\pm 1.05$
5 2.98 $\pm$ 0.43 3.70 $\pm$ 0.61
10 $3.24\pm 0.65$ $4.08\pm 0.82$
20 $3.30\pm 0.57$ $4.19\pm 0.74$
C.7 Visualization on Affordance
In this subsection, we present visualizations of open-world affordances on seen and unseen objects in Fig. A1.
Figure A1: Visualization on Affordance
C.8 Qualitative results Compare with SOTA
@@
C.9 Statistically Insignificant
For the ablation study, we performed paired two‑sample t‑tests, repeating each test five times and reporting the mean p‑value. As summarized in Table A11, the proposed method is significant at the 95% confidence level for the majority of metrics(P-value $\leq$0.05).
Table A11: Two-test Statistically Insignificant
Metric w/o Affordance w/o CFG w/o $l_{\text{penetration}}$ w/o $l_{\text{aff}}$
MPJPE (Seen) $3.1\times 10^{-6}$ $2.0\times 10^{-4}$ $7.5\times 10^{-5}$ $2.6\times 10^{-2}$
MPJPE (Unseen) $4.4\times 10^{-5}$ $1.2\times 10^{-4}$ $5.5\times 10^{-4}$ $2.6\times 10^{-4}$
FOL (Seen) $1.7\times 10^{-3}$ $6.8\times 10^{-5}$ $1.7\times 10^{-5}$ $3.1\times 10^{-5}$
FOL (Unseen) $2.2\times 10^{-2}$ $2.0\times 10^{-3}$ $4.9\times 10^{-2}$ $3.0\times 10^{-1}$
FID (Seen) $2.4\times 10^{-5}$ $1.5\times 10^{-3}$ $1.0\times 10^{-2}$ $2.0\times 10^{-3}$
FID (Unseen) $2.2\times 10^{-5}$ $8.7\times 10^{-5}$ $2.0\times 10^{-4}$ $2.5\times 10^{-4}$
Diversity (Seen) $1.5\times 10^{-1}$ $1.0\times 10^{-1}$ $4.4\times 10^{-2}$ $1.4\times 10^{-1}$
Diversity (Unseen) $1.8\times 10^{-3}$ $3.0\times 10^{-4}$ $4.6\times 10^{-2}$ $4.8\times 10^{-2}$
MModality (Seen) $2.6\times 10^{-4}$ $2.5\times 10^{-2}$ $5.0\times 10^{-3}$ $1.1\times 10^{-2}$
MModality (Unseen) $1.0\times 10^{-3}$ $4.7\times 10^{-2}$ $2\times 10^{-2} 2.2$$2\times 10^{-2}$
@@ Appendix D Code and Dataset
D.1 Code
We will release our code as soon as possible. GitHub is OpenHOI
D.2 Dataset
@@
This dataset contains 571,645 synchronized multi‑view RGB‑D frames captured with five Kinect sensors in three indoor scenes. Each frame includes 3D poses for both hands, 6‑DoF object poses, and verb–noun action labels (36 classes). Split into training (344,645), validation (73,380), and test (153,620) frames, H2O supports egocentric interaction recognition and manipulation benchmarks.
@@
Appendix E Discussion
E.1 Generalizability ability in HOI: Affordance as a key
Affordance is a powerful and explicit prior for interaction that can guide complex and fine-grained HOI synthesis. Our model achieves strong generalization by using affordance as a middleware layer.
  • Open-World Affordance Grounding: We first employ a coarse-to-fine tuning strategy to equip the model with strong affordance reasoning capabilities, enabling it to generate open-world affordance grounding. This enables the synthesis of realistic HOI sequences, demonstrating strong template-free generalization capabilities.
  • Affordance serves as a crucial condition: The open-world affordance grounding serves as a crucial condition for the affordance-driven HOI Diffusion.We incorporate affordance not only during training but also in the loss-guidance applied during inference.
  • Affordance-based Refinement: we design an improved refinement strategy based on affordance: for the interaction between the hand and the object, we optimize based on affordance grounding rather than the conventional closest-surface-point approach.
  • A Template-free Example: Our 3D MLLM has learned from a wide variety of cups, it can still generate accurate affordance grounding for a completely unseen mug. Accurate affordance grounding will guide the synthesis of realistic HOI sequences.
E.2 How to use 3D MLLM in Emboided AI: Choose Powerful Foundation Model and Coarse-to-fine tuning
The 3D multimodal large model has been widely applied in HOI and embodied intelligence, and it is very important to choose a basic model suitable for downstream tasks. In OpenHOI, we chose ShapeLLM as our base model, ShapeLLM is a powerful 3D foundation model that performs exceptionally well on multiple downstream tasks (e.g., Embodied Visual Grounding, Visual Question Answering, and Scene Understanding), making it highly suitable for HOI tasks.After research, we found that ShapeLLM has the following advantages and disadvantages
Advantages: ShapeLLM is trained on a large amount of 3D embodied interaction data and achieves state-of-the-art performance across various downstream tasks. It possesses strong priors in 3D interaction and demonstrates impressive zero-shot 3D representation capabilities.
Disadvantages: ShapeLLM has not been trained on part‑level object annotations, which limits its reasoning capabilities for fine‑grained object understanding.
In order to make the selected 3D base model as suitable as possible for our task, we need to use data to fine tune the model. We adopt a coarse-to-fine tuning strategy: we first pre-train the model on an object-centric affordance dataset to enable it to acquire strong affordance priors. Then, we fine-tune the model on HOI datasets to better align the semantics with the target domain. This allows the model to learn highly effective affordance representations.
E.3 Future of Work: Real-World Applications, about AR/VR and Robotics
OpenHOI can be extended to a wide range of future work in other fields. We have listed several noteworthy areas and provided preliminary solutions for the challenges in future applications
Robotics Manipularion: OpenHOI can be integrated into real‑world robotic manipulation systems, including industrial robot arms and service robots, to enable more flexible and human‑like interactions[wu2024afforddp, zhang2024artigraspphysicallyplausiblesynthesis, zhang2024graspxlgeneratinggraspingmotions].
  • Open-World Affordance Grounding as Powerful Guidance for Robots: Leveraging OpenHOI’s 3D MLLM, our method performs open-world affordance grounding to facilitate the identification of feasible grasping, pushing, and tool-use regions on novel objects, thereby significantly improving success rates for pick-and-place, assembly, and tool-handling tasks.
  • Realistic HOI sequences synthesis for Robot Manipulation: The HOI sequences generated by OpenHOI can be adapted into robotic manipulation sequences. First, we can use an extraction algorithm to obtain the object’s 6-DoF pose. Then, inverse kinematics are applied to the wrist parameters to compute the robot arm’s pose. Finally, a retargeting algorithm transfers the human hand motions onto various robotic hand configurations for manipulation. This approach ensures smooth, precise, and robust manipulation behaviors in real‑world deployments.
Virtual Reality Vision: By synthesizing realistic 3D hand–object interaction sequences, OpenHOI enables users to manipulate virtual objects naturally, for example, by picking up, twisting, or pouring items, thereby enhancing immersion in training simulators, gaming, and virtual prototyping.
Challenges: Robotic manipulation tasks typically demand rapid inference. We plan to employ DPM‑Solver to accelerate the diffusion inference process, which is an important direction for our future work[pan2025unidbfastsamplingunified].